OpenStudy (anonymous):

Write the equation of a line given point (2, 11) and m = 5

OpenStudy (anonymous):

u know the formula for it ?

OpenStudy (anonymous):

plug it in this formula (y-y1)=m(x-x1)

OpenStudy (anonymous):

y =? x +?

OpenStudy (anonymous):

eqn of a line is y = mx + b where y is a point y, x is the point x m is the slop adn b is the y-intercept

OpenStudy (anonymous):

y=y and x=x

OpenStudy (anonymous):

and so since you are given the slope, and a point, plug those into the equation to solve for b.

OpenStudy (anonymous):

you need an equation not an answer as a number

hero (hero):

The equation of the line is y = mx + b You are given the slope m = 5 and one of the points (x,y) = (2,11). Plug everything in then solve for b. 11 = 5(2) + b

OpenStudy (anonymous):

and then you get an equation of y = m(given)x + b(calculated)

OpenStudy (anonymous):

y=mx+b is point slope form of a line. The equation of the line you are looking for is y-y1=m(x-x1). Plug in the given point in y1 and x1

OpenStudy (anonymous):

ya i was telling u the same @EmoLover3 we won't use y=mx+b formula we will user y-y1=m(x-x1)

OpenStudy (anonymous):

so the equation of the line will be y-11=5(x-2)

OpenStudy (anonymous):

y=5x+1 is the equation of the line for your question.

hero (hero):

They're all the equation of the line, but in different forms.

hero (hero):

ax + by = c is the equation of the line in standard form y = mx + b is the equation of the line in slope-intercept form y - y1 = m(x - x1) is the equation of the line in point-slope form

OpenStudy (anonymous):

hmmthe answer will be the same ...... but y-y1=m(x-x1) is easy for now as u just have to plug the values in and get the answer whereas in y=mx+b first u gotta find out b then again plug in the same formula its a little lenghty